OpenStudy (anonymous):

write 1.4x10^3 as a number

OpenStudy (jdoe0001):

\(\large 1.4\times 10^3\qquad \begin{cases} 10^{\underline{3}}\to 1\underline{000}\\ 1.4\times 10^3\to 1.4\times 1\underline{000}\to \square ? \end{cases}\)

OpenStudy (anonymous):

1.4 0 0*1000= -----> 1 4 0 0 ----------------- 4 3 2 1 Look at how the decimal point shift

OpenStudy (anonymous):

1,400